OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? My answer is -8+4x Am I right? :)

OpenStudy (uri):

Where is the equation?

OpenStudy (anonymous):

4x − y = 8 2x + 3y = 9 Sorry about that :)

OpenStudy (anonymous):

Heres the question For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9 8 − 4x 8 + 4x −8 + 4x −8 − 4x

OpenStudy (uri):

4x-8

OpenStudy (uri):

You are correct.

OpenStudy (uri):

4x-8 =-8+4x

OpenStudy (anonymous):

Ok thank you! :)